OpenStudy (anonymous):

-15 - (-3) + 4 simplify A. -16 B. -8 C. -22 D. 8

OpenStudy (anonymous):

@heavenfun13

OpenStudy (midhun.madhu1987):

-15 - ( - 3) + 4 = -15 + 3 + 4 =?

OpenStudy (midhun.madhu1987):

- * - = +

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

-22